# LLM-enhanced Reranking in Recommender Systems

### v2

@@
Abstract
Reranking is a critical component in recommender systems, playing an essential role in refining the output of recommendation algorithms. Traditional reranking models have focused predominantly on accuracy, but modern applications demand consideration of additional criteria such as diversity and fairness. Existing reranking approaches often fail to harmonize these diverse criteria effectively at the model level. Moreover, these models frequently encounter challenges with scalability and personalization due to their complexity and the varying significance of different reranking criteria in diverse scenarios. In response, we introduce a comprehensive reranking framework enhanced by LLM, designed to seamlessly integrate various reranking criteria while maintaining scalability and facilitating personalized recommendations. This framework employs a fully connected graph structure, allowing the LLM to simultaneously consider multiple aspects such as accuracy, diversity, and fairness through a coherent Chain-of-Thought (CoT) process. A customizable input mechanism is also integrated, enabling the tuning of the language model’s focus to meet specific reranking needs. We validate our approach using three popular public datasets, where our framework demonstrates superior performance over existing state-of-the-art reranking models in balancing multiple criteria.
Introduction
@@
Reranking is a fundamental technique in the field of recommender systems (Liu et al. 2022; Zhang et al. 2019; Qin et al. 2022). Its importance lies in its ability to refine and enhance the results generated by ranking models, ultimately providing users with the most relevant and personalized recommendations. In the typical recommendation process, interaction features (e.g., user attributes and item properties) are initially utilized by a ranking model to generate a candidate list for the user. Subsequently, a reranking model is applied to further scrutinize the association between candidate items, delving into the nuances and intricacies of user preferences, and ultimately producing the final recommendation list. Currently, existing reranking models predominantly focus on improving the accuracy aspect of recommendations (Ai et al. 2018; Pei et al. 2019; Lin et al. 2022). Although accuracy is crucial, it is equally important to consider extensive aspects of the recommended list in practical use, such as diversity (Carbonell and Goldstein 1998; Chen, Zhang, and Zhou 2018; Yan et al. 2021) and fairness (Wu et al. 2021; Zehlike et al. 2017). Diversity ensures that users are exposed to a varied range of items, while fairness guarantees equal representation and exposure of different categories or sellers. Though several research (Carbonell and Goldstein 1998; Chen, Zhang, and Zhou 2018; Wu et al. 2021) try to combine one of them with the accuracy aspect for modeling, how to better consider and balance more aspects simultaneously remains a problem.
@@
Specifically, existing reranking models do suffer from several limitations. Firstly, it is challenging to comprehensively consider and balance the complex combination of multiple aspects in modeling, primarily due to the substantial semantic gap between these aspects (Fan et al. 2023; Wang et al. 2022a). This is because each aspect scrutinizes the recommendation list via unique attribute dimensions, highlighting intricate semantic relationships and distinctions. This complexity underscores the substantial gap that exists between different aspects. Furthermore, scalability issues present another major hurdle, inhibiting the application of a singular model across diverse recommendation settings that may prioritize different aspects or functional rules. This challenge is particularly pronounced when introducing novel aspects or custom reranking rules, such as backward rules or stop conditions, not initially anticipated during the model’s development. Moreover, the inability to personalize the amalgamation of various aspects further limits the personalization of existing models, as noted in prior research (Pitoura, Stefanidis, and Koutrika 2022; Zhang and Hurley 2008). Once deployed, the output tendency of a specific model on different aspects is fixed and cannot be intelligently adjusted according to evolving business or user preferences.
@@
Currently, with the rapid development of Large Language Models (LLMs) (Min et al. 2023; Floridi and Chiriatti 2020), extensive research has been undertaken to evaluate the capabilities of LLMs in various contexts (Chang et al. 2023; Wang et al. 2023). Previous studies (Kocoń et al. 2023; Sun et al. 2023; Ma et al. 2023) highlight that zero-shot LLMs may not achieve the same level of proficiency as specialized models in tasks such as information extraction and recommendation. This limitation is often attributed to their restricted token count and difficulties in processing extensive contexts that include thousands of items (Liu et al. 2023). Despite these challenges, these studies have shown that LLMs can perform comparably or even surpass supervised benchmarks in reranking tasks. This efficacy is credited to their robust semantic understanding capabilities within concise contexts involving a limited number of items. Thus, the application of LLMs in the reranking phase of recommendations could serve as a key strategy to merge different aspects by enhancing semantic understanding.
@@
To address these challenges, we propose LLM4Rerank, an innovative reranking framework that harnesses the power of zero-shot LLMs for more precise reranking. Specifically, LLM4Rerank represents various aspect requirements in reranking as distinct nodes, allowing the framework to automatically incorporate these nodes in a Chain-of-Thought (CoT) manner (Wei et al. 2022; Yao et al. 2023; Besta et al. 2023). The advantage of this approach is twofold: it ensures scalability, allowing for the seamless inclusion of new nodes to address emerging aspect requirements. To demonstrate this, in addition to the accuracy aspect, diversity, and fairness aspects are also added for LLM4Rerank modeling in this paper. Additionally, the LLM used in LLM4Rerank can automatically determine the next node to consider, guided by the current reranking history and an additional sentence input referred to as the “Goal”, which is provided by the user or deployer, representing the overall focus and objective of the ongoing reranking process. This dynamic process enables LLM4Rerank to achieve enhanced personalization in the reranking process.
@@
Problem Formulation
The reranking task plays a pivotal role in recommender systems. As depicted in Figure 1, consider as the set of users and as the set of items available for recommendation. For clarity, this paper represents each user and item by a feature vector ( and , respectively). Initially, a ranking model generates a candidate item list with items for each user. To improve recommendation performance, a reranking process (Pei et al. 2019; Liu et al. 2022) is applied to analyze the relationships among items within the initial list . This analysis aims to generate a refined list with items, denoted as , from the initial list. The goal of reranking models is to enhance user-item relevance by optimizing a defined objective function:
@@
In this study, to facilitate equitable comparisons across various reranking baselines, the Generalized Matrix Factorization (GMF) model (He et al. 2017) is employed as the uniform global ranking model following previous study (Lin et al. 2022). This approach ensures that all reranking models operate on identical candidate item lists.
LLM4Rerank Overview
@@
Nodes Construction
To facilitate the Large Language Model (LLM)’s systematic analysis of complex aspect requirements in reranking, this structure aims to establish distinct nodes for specific requirements. Such an arrangement enables the LLM to process these requirements in a Chain-of-Thought approach (Wei et al. 2022; Yao et al. 2023; Besta et al. 2023). Nevertheless, this approach presents two primary challenges: First, customizing the node structure to maintain scalability when incorporating additional requirements; and second, empowering the LLM to automatically select its subsequent reranking step. To address these challenges, we introduce a generic node structure. It comprises a reranking step, paired with an ancillary indicator that signifies the direction of the forthcoming step identified by the next node’s name. This configuration permits the LLM to automatically navigate through the LLM4Rerank framework, making decisions based on the presently available information.
This section outlines our strategy for addressing the challenge of node structure customization with the aim of augmenting the scalability of the LLM4Rerank framework. Specifically, we introduce a generic node structure that serves as the foundation for all nodes as shown in Figure 2 (b). This generic node represents a single step of reranking under LLM considerations. The inputs to a generic node encompass semantic representations of user information, candidate items, the “Goal” sentence that defines the personalized focus for the entire reranking process, and the whole historical reranking pool if available. Outputs from this node are twofold: the immediate reranking results for the current node, represented by a list of item IDs, are integrated into the historical reranking pool alongside the current node’s name, serving as a reference for subsequent steps. Additionally, an indicator (i.e., the next node’s name in this paper) specifying the subsequent node for reranking is produced, thereby achieving an automatic step-by-step procedure. Within each node, LLM4Rerank initially crafts a prompt tailored to the specific reranking criteria and inputs at hand with a predefined template. Subsequently, LLM4Rerank would interact with LLM to obtain the two outputs based on the generated prompt.
Aspect Nodes
To facilitate the LLM in executing reranking tasks tailored to distinct aspect requirements, we employ a prompt-based template approach within the proposed generic node structure. This method allows for the instantiation of specific nodes dedicated to evaluating different aspects within the reranking process. Consequently, each node is designed to systematically address one of these key aspects, ensuring that the reranking outcomes reflect a balanced consideration. In this study, to demonstrate the scalability of LLM4Rerank, we implement three aspect nodes dedicated to reranking: “Accuracy”, “Diversity”, and “Fairness”.
@@
Figure 3: Example prompt template of the accuracy node.
Figure 4: Example prompt template of the diversity node. -
Diversity Node: This node is specifically designed to address the diversity criteria for the final recommendation list during the reranking phase. In this study, we assess the diversity of the reranking outcomes by evaluating the extent to which a particular attribute of the items is varied within the final list. We employ the -NDCG metric (Clarke et al. 2008) for this purpose. Consequently, an illustrative example of the template used in the diversity node is depicted in Figure 4.
Fairness Node: This node is designated to meet the fairness objectives within the final recommendation list at the reranking phase. In our study, fairness of the recommendation outcomes is operationalized as the average score disparity across two sample groups, segregated by a distinct characteristic, and evaluated using the Mean Absolute Deviation (MAD) metric (Zhu, Hu, and Caverlee 2018). Given that the LLM inherently generates reranking lists rather than numerical scores, we allocate scores ranging linearly from 1 to 0 to the items in the final recommendation list. These scores are subsequently utilized to compute the MAD for fairness assessment. For an in-depth methodological exposition, readers are directed to Section Implementation Details. Figure 5 provides a straightforward template illustration for the fairness node.
Figure 5: Example prompt template of the fairness node.
Functional Nodes
Recent research has demonstrated the efficacy of reflection in optimizing the output of LLMs (Shinn, Labash, and Gopinath 2023; Ji et al. 2023). To augment the logical capabilities of LLM4Rerank in the reranking process and introduce specialized functionalities, we have developed two functional nodes specifically aimed at facilitating reflection and termination within the reranking sequence.
@@
Automatic Reranking Process
@@
Automatic Transition Across Nodes: For each node, upon receiving replies from LLM, LLM4Rerank would obtain the current reranking results, along with an indicator (i.e., the next node name) for the subsequent node, thereby ensuring a fluid and automatic transition across various nodes.
@@
By applying these sub-processes, the whole automatic reranking process of LLM4Rerank is established as illustrated in Algorithm 1.
Output: Final reranking result
Note: represents the execution of functions in node with input .
Experiments
@@
Experimental Setup
Dataset
We conduct our experiments using three widely recognized public datasets: ML-1M 111https://grouplens.org/datasets/movielens/1m/ (Harper and Konstan 2015), KuaiRand (KuaiRand-Pure) 222https://kuairand.com/ (Gao et al. 2022), and Douban-Movie 333https://www.kaggle.com/datasets/fengzhujoey/douban-datasetratingreviewside-information (Zhu et al. 2020, 2019). For each dataset, we employ the leave-one-out method (He et al. 2017; Bayer et al. 2017; Gan et al. 2021), a widely adopted approach in the literature, for dividing the data into training, validation, and testing sets. According to previous studies (Lin et al. 2022), we select the Generalized Matrix Factorization (GMF) model as the global ranking model to generate candidate lists for each user, with each set comprising 20 items. To ensure equitable comparisons across deep learning and LLM-based models, we omit features lacking explicit semantic information and exclude users with fewer than five interactions. For deep learning-based models, we utilize a standard embedding technique (Wang et al. 2021; Guo et al. 2021) to transform various features into vector inputs. Conversely, for LLM-based models, the semantic information of the feature (e.g., the feature’s name) is utilized as the input. Table 1 presents the statistics of both datasets subsequent to preprocessing.

### v3

# LLM4Rerank: LLM-based Auto-Reranking Framework for Recommendations

@@
Abstract.
Reranking is significant for recommender systems due to its pivotal role in refining recommendation results. Numerous reranking models have emerged to meet diverse reranking requirements in practical applications, which not only prioritize accuracy but also consider additional aspects such as diversity and fairness. However, most of the existing models struggle to strike a harmonious balance between these diverse aspects at the model level. Additionally, the scalability and personalization of these models are often limited by their complexity and a lack of attention to the varying importance of different aspects in diverse reranking scenarios. To address these issues, we propose LLM4Rerank, a comprehensive LLM-based reranking framework designed to bridge the gap between various reranking aspects while ensuring scalability and personalized performance. Specifically, we abstract different aspects into distinct nodes and construct a fully connected graph for LLM to automatically consider aspects like accuracy, diversity, fairness, and more, all in a coherent Chain-of-Thought (CoT) process. To further enhance personalization during reranking, we facilitate a customizable input mechanism that allows fine-tuning of LLM’s focus on different aspects according to specific reranking needs. Experimental results on three widely used public datasets demonstrate that LLM4Rerank outperforms existing state-of-the-art reranking models across multiple aspects.
Relevance Statement: This paper presents an LLM-based auto-reranking framework that integrates various personalized aspects during the reranking stage of a recommender system. This is highly pertinent to the Web track topic user modeling, personalization, and recommendation. Moreover, the proposed method can be directly used to conduct reranking with personalized requirements, which is closely related to the applications of the Web.
1. Introduction
@@
Reranking is a fundamental technique in the field of recommender systems (Liu et al., 2022b; Gao et al., 2024; Li et al., 2023b). Its importance lies in its ability to refine and enhance the results generated by ranking models, ultimately providing users with the most relevant and personalized recommendations. In the typical recommendation process, interaction features (e.g., user attributes and item properties) are initially utilized by a ranking model to generate a candidate list for the user. Subsequently, a reranking model is applied to further scrutinize the association between candidate items, delving into the nuances and intricacies of user preferences, and ultimately producing the final recommendation list. Currently, existing reranking models predominantly focus on improving the accuracy aspect of recommendations (Ai et al., 2018; Pei et al., 2019; Lin et al., 2022; Zhao et al., 2018a). Although accuracy is crucial, it is equally important to consider extensive aspects of the recommended list in practical use, such as diversity (Carbonell and Goldstein, 1998; Chen et al., 2018; Yan et al., 2021) and fairness (Wu et al., 2021; Zehlike et al., 2017). Diversity ensures that users are exposed to a varied range of items, while fairness guarantees equal representation and exposure of different categories or sellers. Though several research (Carbonell and Goldstein, 1998; Chen et al., 2018; Wu et al., 2021) try to combine one of them with the accuracy aspect for modeling, how to better consider and balance more aspects simultaneously remains a problem.
@@
Specifically, existing reranking models do suffer from several limitations. Firstly, it is challenging to comprehensively consider and balance the complex combination of multiple aspects in modeling, primarily due to the substantial semantic gap between these aspects (Fan et al., 2023; Wang et al., 2022b). This is because each aspect scrutinizes the recommendation list (Liu et al., 2023a, 2022a; Zhao et al., 2023; Liu et al., 2024a) via unique attribute dimensions, highlighting intricate semantic relationships and distinctions (Xu et al., 2024b). This complexity underscores the substantial gap that exists between different aspects. Furthermore, scalability issues present another major hurdle, inhibiting the application of a singular model across diverse recommendation settings that may prioritize different aspects or functional rules. This challenge is particularly pronounced when introducing novel aspects or custom reranking rules, such as backward rules or stop conditions, not initially anticipated during the model’s development. Moreover, the inability to personalize the amalgamation of various aspects further limits the personalization of existing models, as noted in prior research (Pitoura et al., 2022). Once deployed, the output tendency of a specific model on different aspects is fixed and cannot be intelligently adjusted according to evolving business or user preferences.
@@
Currently, with the rapid development of Large Language Models (LLMs) (Min et al., 2023; Floridi and Chiriatti, 2020; Jia et al., 2024; Liu et al., 2024c; Fu et al., 2023), extensive research has been undertaken to evaluate the capabilities of LLMs in various contexts (Chang et al., 2023; Wang et al., 2023b; Li et al., 2023a; Zhang et al., 2024). Previous studies (Kocoń et al., 2023; Sun et al., 2023; Ma et al., 2023) highlight that zero-shot LLMs may not achieve the same level of proficiency as specialized models in tasks such as information extraction and recommendation. This limitation is often attributed to their restricted token count and difficulties in processing extensive contexts that include thousands of items (Liu et al., 2023b). Despite these challenges, these studies have shown that LLMs can perform comparably or even surpass supervised benchmarks in reranking tasks. This efficacy is credited to their robust semantic understanding capabilities within concise contexts involving a limited number of items. Thus, the application of LLMs in the reranking phase of recommendations could serve as a key strategy to merge different aspects by enhancing semantic understanding.
@@
To address these challenges, we propose LLM4Rerank, an innovative reranking framework that harnesses the power of zero-shot LLMs for more precise reranking. Specifically, LLM4Rerank represents various aspect requirements in reranking as distinct nodes, allowing the framework to automatically incorporate these nodes in a Chain-of-Thought (CoT) manner (Wei et al., 2022; Yao et al., 2023; Besta et al., 2023). The advantage of this approach is twofold: it ensures scalability, allowing for the seamless inclusion of new nodes to address emerging aspect requirements. To demonstrate this, in addition to the accuracy aspect, diversity, and fairness aspects are also added for LLM4Rerank modeling in this paper. Additionally, the LLM used in LLM4Rerank can automatically determine the next node to consider, guided by the current reranking history and an additional sentence input referred to as the “Goal”, which is provided by the user or deployer, representing the overall focus and objective of the ongoing reranking process. This dynamic process enables LLM4Rerank to achieve enhanced personalization in the reranking process.
@@ 2. Framework
2.1. Problem Formulation
The reranking task plays a pivotal role in recommender systems. As depicted in Figure 1, consider as the set of users and as the set of items available for recommendation. For clarity, this paper represents each user and item by a feature vector ( and , respectively). Initially, a ranking model generates a candidate item list with items for each user. To improve recommendation performance, a reranking process (Pei et al., 2019; Liu et al., 2022b) is applied to analyze the relationships among items within the initial list . This analysis aims to generate a refined list with items, denoted as , from the initial list. The goal of reranking models is to enhance user-item relevance by optimizing a defined objective function:
@@ 2.1. Problem Formulation
In this study, to facilitate equitable comparisons across various reranking baselines, the Generalized Matrix Factorization (GMF) model (He et al., 2017) is employed as the uniform global ranking model following previous study (Lin et al., 2022). This approach ensures that all reranking models operate on identical candidate item lists.
2.2. LLM4Rerank Overview
@@
2.3. Nodes Construction
To facilitate the Large Language Model (LLM)’s systematic analysis of complex aspect requirements in reranking, this structure aims to establish distinct nodes for specific requirements. Such an arrangement enables the LLM to process these requirements in a Chain-of-Thought approach (Wei et al., 2022; Yao et al., 2023; Besta et al., 2023). Nevertheless, this approach presents two primary challenges: First, customizing the node structure to maintain scalability when incorporating additional requirements; and second, empowering the LLM to automatically select its subsequent reranking step. To address these challenges, we introduce a generic node structure. It comprises a reranking step, paired with an ancillary indicator that signifies the direction of the forthcoming step identified by the next node’s name. This configuration permits the LLM to automatically navigate through the LLM4Rerank framework, making decisions based on the presently available information.
This section presents our strategy to address the challenge of node structure customization and enhance the scalability of the LLM4Rerank framework by introducing a customizable, generic node structure, as depicted in Figure 2 (b). Each node performs a reranking step under LLM considerations, defined by the equation:
Inputs for each node include user information , candidate items , a personalized sentence for illustrating personalized focus for the entire reranking process, and the historical reranking pool , if available. Outputs consist of reranked result and the next node’s name which indicates further reranking steps. The LLM4Rerank framework crafts prompt specific to reranking criteria and interacts with LLM to generate these outputs.
2.3.1. Aspect Nodes
To facilitate the LLM in executing reranking tasks tailored to distinct aspect requirements, we employ a prompt-based template approach within the generic node structure:
where is the prompt template for different nodes. This method allows for instantiating specific nodes dedicated to evaluating different aspects within the reranking process. Consequently, each node is designed to systematically address one of these key aspects, ensuring that the reranking outcomes reflect a balanced consideration. In this study, to demonstrate the scalability of LLM4Rerank, we implement three aspect nodes dedicated to reranking: “Accuracy”, “Diversity”, and “Fairness”. Note that the example prompts below are simplified. In practice, detailed prompts should be tailored to input features, aspect characteristics, and evaluation metric semantics.
@@ 2.3.1. Aspect Nodes
Figure 3. Example prompt template of the accuracy node. -
Diversity Node: This node is specifically designed to address the diversity criteria for the final recommendation list during the reranking phase. In this study, we assess the diversity of the reranking outcomes by evaluating the extent to which a particular attribute of the items is varied within the final list. We employ the -NDCG metric (Clarke et al., 2008) for this purpose. Consequently, an illustrative example of the template used in the diversity node is depicted in Figure 4.
Figure 4. Example prompt template of the diversity node. -
Fairness Node: This node is designated to meet the fairness objectives within the final recommendation list at the reranking phase. In our study, fairness of the recommendation outcomes is operationalized as the average score disparity across two sample groups, segregated by a distinct characteristic, and evaluated using the Mean Absolute Deviation (MAD) metric (Zhu et al., 2018). Given that the LLM inherently generates reranking lists rather than numerical scores, we allocate scores ranging linearly from 1 to 0 to the items in the final recommendation list. These scores are subsequently utilized to compute the MAD for fairness assessment. For an in-depth methodological exposition, readers are directed to Section 3.1.3. Figure 5 provides a straightforward template illustration for the fairness node.
Figure 5. Example prompt template of the fairness node.
2.3.2. Functional Nodes
Recent research has demonstrated the efficacy of reflection in optimizing the output of LLMs (Shinn et al., 2023; Ji et al., 2023). To augment the logical capabilities of LLM4Rerank in the reranking process and introduce specialized functionalities, we have developed two functional nodes specifically aimed at facilitating reflection and termination within the reranking sequence.
@@ 2.3.2. Functional Nodes
Figure 6. Example prompt template of the backward node. -
@@
2.4. Automatic Reranking Process
@@
Automatic Transition Across Nodes: For each node, upon receiving replies from LLM, LLM4Rerank would obtain the current reranking results, along with an indicator (i.e., the next node name) for the subsequent node, thereby ensuring a fluid and automatic transition across nodes as illustrated in Equation (2).
@@
By applying these sub-processes, the whole automatic reranking process of LLM4Rerank is established. The overall Algorithm is also provided in Appendix A.
3. Experiments
@@
3.1. Experimental Setup
3.1.1. Dataset
We conduct our experiments using three widely recognized public datasets: ML-1M 111https://grouplens.org/datasets/movielens/1m/ (Harper and Konstan, 2015), KuaiRand (KuaiRand-Pure) 222https://kuairand.com/ (Gao et al., 2022), and Douban-Movie 333https://www.kaggle.com/datasets/fengzhujoey/douban-datasetratingreviewside-information (Zhu et al., 2020, 2019). For each dataset, we employ the leave-one-out method (He et al., 2017; Bayer et al., 2017; Gan et al., 2021), a widely adopted approach in the literature, for dividing the data into training, validation, and testing sets. According to previous studies (Lin et al., 2022), we select the Generalized Matrix Factorization (GMF) model as the global ranking model to generate candidate lists for each user, with each set comprising 20 items. To ensure equitable comparisons across deep learning and LLM-based models, we omit features lacking explicit semantic information and exclude users with fewer than five interactions. For deep learning-based models, we utilize a standard embedding technique (Wang et al., 2021; Guo et al., 2021) to transform various features into vector inputs. Conversely, for LLM-based models, the semantic information of the feature (e.g., the feature’s name) is utilized as the input. Table 1 presents the statistics of both datasets subsequent to preprocessing.
3.1.2. Baseline
@@
GMF (He et al., 2017) extends matrix factorization into a non-linear framework, serving as the primary global ranking method in this study. The GMF results represent recommendations prior to the application of any reranking process.
DLCM (Ai et al., 2018) enhances reranking efficacy by employing a recurrent neural network alongside an attention-based loss function to comprehend local ranking dynamics, aiming primarily at improving accuracy within recommendation outcomes.
PRM (Pei et al., 2019) leverages a transformer architecture with self-attention mechanisms to refine the entire recommendation list by acknowledging the inter-item influences, thereby concentrating on augmenting accuracy.
MMR (Carbonell and Goldstein, 1998) aims to balance query relevance with the reduction of redundancy in reranked documents, employing a maximal marginal relevance score to bolster the diversity aspect in recommendation outcomes.
FastDPP (Chen et al., 2018) expedites Determinantal Point Processes (DPP) for Maximum A Posteriori (MAP) inference, facilitating the efficient production of diverse recommendation sets. This model focuses on the diversity aspect of the recommendation result.
FairRec (Wu et al., 2021) introduces a fairness-aware recommendation framework that employs decomposed adversarial learning and orthogonality regularization. It aims to alleviate bias concerning sensitive user attributes, thereby fostering fairness in recommendations without compromising overall performance.
RankGPT (Sun et al., 2023) investigates the application of LLMs in ranking tasks within information retrieval, employing a novel instructional permutation generation method alongside a sliding window strategy. This model is distinguished by its focus on accuracy. Note that, as a zero-shot LLM baseline, the permutation distillation method in the original paper is not implemented.
GoT (Besta et al., 2023) proposes a graph-of-thought approach to enhance LLMs’ prompting efficacy by structuring generated content as a graph. This facilitates synergistic outcomes, thought distillation, and feedback loop integration, aligning LLM reasoning more closely with human cognitive processes. Unlike LLM4Rerank, GoT adheres to predetermined node-to-node inference paths without historical data consideration. By applying a fixed path, “Accuracy-Diversity-Fairness-Stop.” GoT serves as a zero-shot LLM baseline focusing on the combination of accuracy, diversity, and fairness aspects in this paper.
3.1.3. Implementation Details
In the evaluation of the accuracy aspect, we adopt widely recognized metrics: Hit Ratio (HR) and Normalized Discounted Cumulative Gain (NDCG). For assessing the diversity aspect, we apply the commonly used metric -NDCG. To evaluate fairness, we use the Mean Absolute Difference (MAD) (Zhu et al., 2018). The MAD calculation is formalized as:
where and represent the predicted ratings for two distinct groups, and denotes the total number of ratings for group . Within the ML-1M dataset, we utilize the “genre” feature for diversity analysis and the “year” feature for fairness assessment, categorizing films into two groups based on their release year: pre-1996 and post-1996 (Zhu et al., 2018; Kamishima and Akaho, 2017). For the KuaiRand dataset, “upload_type” serves as the criterion for diversity, while “video_duration”-divided into less than 60,000 ms and greater than 60,000 ms-serves for fairness evaluation. To fine-tune deep learning-based models for optimal performance, we set the learning rate at 0.001 and engage in a grid search to determine the best hyper-parameters. For zero-shot LLM baselines and LLM4Rerank, Llama-2-13B (Touvron et al., 2023) is selected as the default LLM backbone. In addition, we also provide hyper-parameter analysis, inference analysis, and reproduction guideline in the Appendix B, C, D, and E.